OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find 2·f(4).

OpenStudy (johnweldon1993):

Well we are given g(x) but do not need it here. It looks like f(x) = x^2 + 1 correct? So f(4) would be that but replacing all the 'x' with 4...so f(4) = (4)^2 + 1 = 16 + 1 = 17 so f(4) = 17 And 2 * f(4) would be 2 * 17 = 34